OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 2, -4, and 1 + 3i

OpenStudy (anonymous):

@SolomonZelman @phi

OpenStudy (anonymous):

@ParthKohli

OpenStudy (anonymous):

From the given roots, the polynomial can be factored to look like this: \[(x-2)(x+4)(x-(1+3i))\] As is, if you were to expand it, you'd have some complex coefficients. To rectify that, recall that complex roots come in conjugate pairs, so \(1-3i\) is also a root, and you have \[(x-2)(x+4)(x-(1+3i))(x-(1-3i))\] Expand and write in standard form.

OpenStudy (anonymous):

@SithsAndGiggles how would I write it in standard form?

OpenStudy (anonymous):

Standard form of a polynomial of degree \(n\) is \[a_nx^n+a_{n-1}x^{n-1}+\cdots+a_2x^2+a_1x+a_0\] So for example, if you have \(f(x)=(x-1)(x+1)\), then standard form is obtained by expanding: \[f(x)=x^2-x+x-1=x^2-1\]